# UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities

### v1

@@ 6.4.1. Analysis of Contrastive Learning
Table 8. Comparison experiments with different retrieval contents on retrieval augmentation.
Method Metric Type MAP Avg
@10 @20 @50 @100
RetExpan + RA (Entity Introduction) Pos 44.74 42.23 40.66 40.27 41.97
Neg 5.53 5.88 6.56 7.44 6.35
Comb 69.60 68.18 67.05 66.42 67.81
RetExpan + RA (Wikidata Attributes) Pos 40.94 38.78 37.28 37.64 38.66
Neg 6.93 6.79 7.89 9.56 7.79
Comb 67.00 66.00 64.69 64.04 65.43
RetExpan + RA (GT Attributes) Pos 45.28 43.84 43.77 45.22 44.52
Neg 6.21 6.69 7.92 8.19 7.25
Comb 69.54 68.57 67.92 68.51 68.64
GenExpan + RA (Entity Introduction) Pos 48.68 46.26 43.80 41.27 45.00
Neg 7.16 8.54 8.63 7.70 8.01
Comb 70.76 68.86 67.58 66.78 68.50
GenExpan + RA (Wikidata Attributes) Pos 48.30 45.94 43.91 41.50 44.91
Neg 7.00 8.46 8.79 7.83 8.02
Comb 70.65 68.74 67.56 66.83 68.45
GenExpan + RA (GT Attributes) Pos 49.14 46.83 43.98 42.51 45.62
Neg 6.76 8.43 8.56 7.66 7.85
Comb 71.19 69.20 67.71 67.42 68.88
6.4.2. Analysis of Retrieval Augmentation
In this subsection, we delve into the impact of additional entity knowledge retrieved from various sources on RetExpan and GenExpan, including entity introductions and attribute information extracted from Wikidata, as well as ground-truth positive and negative attribute information. The experiment results presented in Table 8 demonstrate that each of the three knowledge is beneficial for both the original RetExpan and GenExpan.
Concerning the entity attribute information sourced from Wikidata, we randomly sampled an attribute subset and converted it into text with templates. The attributes on Wikidata are intricate and encompass numerous rarely-used attributes (e.g., the attributes of the entity ”Motorola” include its ”YouTube channel ID”). This may lead to the retrieved attributes are not positive or negative attributes that the current ultra-fine-grained semantic class really cares about, thereby hindering model comprehension. Therefore, despite the high quality of Wikidata’s attribute information, its gain to the model is relatively marginal.
Meanwhile, we can easily find that the direct inclusion of ground-truth positive and negative attribute information related to ultra-fine-grained semantic classes has the most substantial gain to the model. Yet, we posit that the potential of retrieval augmentation methods has not been fully realized. Firstly, the supplementary knowledge retrieved for each entity is static across different sentences and does not adapt to the entity’s context, potentially leading to homogenized entity representations. Crafting dynamic and ultra-fine-grained retrieval strategies deserves further exploration. Secondly, simply concatenating retrieved knowledge with the original sentences is not the optimal way to leverage it. Exploring more effective methods for knowledge exploitation is also essential.
6.4.3. Analysis of Chain-of-thought Reasoning
Table 9. Comparison experiments on chain-of-thought reasoning. The impact of reasoning depth and reasoning precision on performance is investigated. GT: Ground-truth. Gen: Generated. CN: Class name. Pos/Neg: Positive/Negative attribute information.
@@ 6.4.3. Analysis of Chain-of-thought Reasoning
Experiments in Table 9 examine the chain-of-thought with varying reasoning depths and assess the impact of reasoning precision on model performance. From the results, we observe that:
(1) Deeper reasoning proves to be more advantageous for Ultra-ESE, as it facilitates a more explicit understanding of the implicit ultra-fine-grained semantics of seed entities.
(2) Surprisingly, the use of manually labeled ground-truth class names is not superior to those derived through reasoning with the LLaMA-7B. Case study of the generated class names reveal that LLaMA-7B can produce class names encapsulating positive attribute information based on the given positive seed entities. For instance, when given entities about U.S. airports with the “location” attribute, GenExpan accurately infers more specific semantic classes that reflect positive attribute information, e.g., “Airports in Michigan”. This suggests that reasoning results are not necessarily inferior to manually annotated “ground-truth” labels, as chain-of-thought reasoning induces LLMs to attend to details that human ignores.
@@
6.5. Explore Experiments on Different Paradigm Model Interactions
Table 10. Experiments on the interaction of RetExpan and GenExpan.
Method Metric Type MAP Avg
@10 @20 @50 @100
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.44
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan + GenExpan Pos 45.93 44.37 42.96 42.35 43.90
Neg 7.03 7.51 9.04 9.48 8.27
Comb 69.45 68.43 66.96 66.43 67.82
GenExpan Pos 46.84 44.99 42.88 40.79 43.88
Neg 7.27 8.29 8.74 8.02 8.08
Comb 69.79 68.35 67.07 66.38 67.90
GenExpan + RetExpan Pos 48.91 46.09 44.13 42.49 45.41
Neg 6.24 6.80 7.56 7.84 7.11
Comb 71.33 69.64 68.29 67.32 69.15
We are curious whether retrieval-based and generation-based frameworks can interact and mutually reinforce each other. Actually, in the process of enhancing the retrieval-based RetExpan, we have leveraged the generative LLMs GPT-4. At the same time, when enhancing the generation-based GenExpan, we have utilized the retrieval-based approach to retrieve extra entity knowledge. In this section, we explore the facilitative relationship between the two paradigms at a higher level. Specifically, for the A+B model in Table 10, we first allow Model A to obtain 1000 entities from the complete candidate entities, which guarantees high entity recall. Then, model B executes on the obtained 1000 entities. To our surprise, no matter whether RetExpan or GenExpan is deployed first, both of them enhance the performance on all metrics considerably. These two different paradigm frameworks have their own advantages. The retrieval-based framework measures the semantic relation between candidate and seed entities by feature similarity, which inherently fits the nature of the retrieval task. Thus only by utilizing smaller models, retrieval-based frameworks can achieve competitive results, and be easily migrated to specific domains. Generation-based frameworks, on the other hand, typically rely on LLMs, which offer a broader knowledge base and greater semantic awareness, but the LLMs tend to be less efficient. Therefore, how to make the two types of models learn collaboratively, promote each other, and then achieve automatic evolution is a promising direction for solving Ultra-ESE.
Appendix A Prompts
When constructing contrastive learning training data, we use GPT’s classification to obtain $L_{pos}$ and $L_{neg}$. For each entity in the top-$T$ of $L_{0}$, determine whether its attributes are consistent with $S^{pos}$ ($S^{neg}$). If it is consistent, GPT should output 1; otherwise, GPT should output 0. Next, these entities that GPT deems consistent with the attributes of $S^{pos}$ ($S^{neg}$) will be merged with $S^{pos}$ ($S^{neg}$) to form $L_{pos}$ ($L_{neg}$). The prompt to classify entity is shown in Table 13.
In GenExpan, each round uses three example entities to construct $Prompt_{g}$, which is used to generate entities that are semantically similar to them. $Prompt_{g}$ is as shown in Table 14.
In the process of enhancing GenExpan with Chain of Thought prompting, we employ LLM to generate fine-grained class names. The prompt, $Prompt_{c}$, is shown in Table 15.
Appendix B Implementation Details of RetExpan
In RetExpan, sentences are tokenized using the WordPiece tokenizer and then fed into a 12-layer Transformer initialized with $BERT_{BASE}$ weights. To optimize training efficiency and preserve semantic knowledge learned by BERT, we freeze the first 11 layers of the encode, thus only the last layer is fine-tuned. When tokenizing, to ensure that mentions of entities in long sentences are not truncated, we have implemented entity focus, guaranteeing the presence of entities within the sentence. During expansion, we use one-shot expansion to directly obtain the preliminary expansion results $L_{0}$. In RetExpan with contrastive learning strategy, each epoch during training alternates between computing and optimizing entity prediction loss and contrastive learning loss.
To ensure that the encoder acquires knowledge from the corpus, we trained it for 20 epochs on 8 RTX 3090 GPUs. The hyperparameters for training learning rate, batch size, weight decay, label smoothing factor $\eta$ were set to 4e-5, 128, 1e-2, and 0.075 respectively.
Appendix C Implementation Details of GenExpan
In GenExpan, we use LLaMA-7b as the base model. Initially, we train for 1 epoch on the corpus using 6 A100 GPUs. The training hyperparameters learning rate, batch size, gradient accumulation steps, weight decay, gradient clipping are set to 1e-5, 4, 8, 1e-4, and 1.0 respectively.
During entity generation in GenExpan, we utilize prefix-constrained beam search with a beam size of 40 to generate 40 entities at a round. For entity selection, we compute the positive similarity score for each entity and select those whose scores are in the top 0.7 as the results for the current round. When no new entity is generated for 20 consecutive rounds, the generation process will end and move to re-ranking.
Appendix D Details of UltraWiki
UltraWiki contains 10 fine-grained semantic classes, which fall under 5 coarse-grained semantic categories: Organization, Location, Product, Person, and Miscellaneous. The number of entities in each fine-grained semantic class ranges from 45 to 952. Each fine-grained semantic class has 2 to 3 attributes, and depending on the combination of attributes, several ultra-fine-grained semantic classes can be derived from a single fine-grained class. Detailed data is displayed in Table 11.
The number of combinations from positive and negative attributes varies significantly, leading to substantial differences in the number of ultra-fine-grained semantic classes. The count of UltraWiki’s positive and negative attributes and their corresponding ultra-fine-grained classes are presented in Table 12.
Figure 7. Parameter analysis experiments.
Table 11. Fine-grained semantic classes detail. CLS.: Semantic Class
Coarse CLS. Fine-grained CLS. #Entities #Ultra-fine-grained CLS. Attributes
Organization Canada universities 99 10 ¡Loc-Province¿, ¡Type¿
Location China cities 675 50 ¡Province¿, ¡Prefecture¿
Countries 190 68 ¡Continent¿, ¡Driving-Side¿, ¡Per-Capita-Income¿
US airports 370 74 ¡Role¿, ¡Loc-State¿
US national monuments 112 12 ¡Loc-State¿ ¡Agency¿
Product Mobile phone brands 159 7 ¡Loc-Continent¿, ¡Status¿
Percussion instruments 128 10 ¡Type¿, ¡Source-Continent¿
Person Nobel laureates 952 11 ¡Prize¿, ¡Gender¿
US presidents 45 5 ¡Party¿, ¡Birth-State¿
Miscellaneous Chemical elements 118 14 ¡Period¿, ¡Phase-at-R.T.¿
Table 12. Types of ultra-fine-grained semantic classes. CLS.: Semantic Class
$|\mathcal{A}^{pos}|$ $|\mathcal{A}^{neg}|$Ultra-fine-grained CLS.
1 1 238
1 2 5
2 1 9
2 2 7
3 3 2
Appendix E Parameter analysis
To explore the sensitivity of each key hyperparameter in RetExpan and GenExpan, we conducted a comprehensive analysis of the smoothing factor $\eta$ and the re-ranking segment length $l$ in RetExpan, as well as the number of positive and negative entities mined in the contrastive learning strategy $|L_{pos}|$ and $|L_{neg}|$, the Top-$p$ of entity selection module in GenExpan, and the re-ranking segment length $l$, respectively. The results are depicted in Figure 7, from which we can observe that:
(1) The label smoothing $\eta$ serves to mitigate the penalty for entities possessing similar semantics to ground-truth entities. In Ultra-ESE, there are fewer entities exhibiting similarity to ground-truth entities at the ultra-fine-grained semantic level, thereby diminishing the utility of label smoothing. Consequently, the fluctuation in model performance due to parameter variations is relatively not significant.
(2) Excessive segment length $l$ in both RetExpan and GenExpan equates to degrading the segmented re-ranking strategy to a naive re-ranking approach, thereby introducing noise from irrelevant entities, as mentioned in the main text. Hence, there is a general decrease in both PosMAP@K as the segment length increases, with larger K values leading to greater declines.
(3) Regarding the ultra-fine-grained contrastive learning strategy in RetExpan, we observe that mining too many or too few positive and negative entities using GPT-4 results in performance degradation. Insufficient positive and negative entities fail to fully stimulate the potential of contrast learning, whereas an excess of such entities introduces noisy entities. Therefore, as the number of entities to be mined increases gradually, PosMAP initially rises and then falls, and NegMAP exhibits a slight initial decline followed by a rise. However, overall, these two parameters demonstrate robustness, with even the worst-performing set of values outperforming the original RetExpan.
(4) A trade-off is also necessary for the Top-$p$. A too small Top-$p$ may fail to ensure the diversity of newly generated entities, while a too large Top-$p$ may introduce an excessive number of irrelevant entities during iterative expansion. Consequently, we observe a decrease in PosMAP and a notable increase in NegMAP when Top-p is increased from 80% to 90%, which indicates that there is considerable space for improving the current rejection strategy for negatively targeted entities.
Appendix F Analysis of Model Size
We utilize various families and sizes of LLMs as backbone model of GenExpan, including 560M, 1B7, 3B, 7B1 for BLOOM, as well as 7B and 13B for LLaMA. The experiment results depicted in Figure 8 align well with the expectation. Both the BLOOM and LLaMA families satisfy the scaling law: larger models tend to be more effective in Ultra-ESE, yielding better results. Moreover, LLaMA-7B outperforms BLOOM-7B1 at the same scale.
Figure 8. Comparison experiments on different LLM families and sizes.
@@ Appendix F Analysis of Model Size
Figure 9. Case studies of GenExpan and the retrieval augmentation, chain-of-thought reasoning techniques. $+$ and - - - stand for positive and negative target entities, respectively. !!! represents irrelevant entities belonging to the same fine-grained semantic class as the seed entities.
Appendix G Case Study
@@
Figure 9 presents some case studies of GenExpan and the corresponding retrieval augmentation, chain-of-thought reasoning techniques. Overall, it is evident that GenExpan predominantly avoids expanding entities that do not belong to the same semantic class as the positive and negative seed entities. This emphasizes the relative tractability of traditional ESE tasks compared to the significant challenges posed by Ultra-ESE for existing LLMs, where the model cannot exclude the expansion of negative target entities and irrelevant entities (belonging to the same fine-grained semantic class).
For retrieval augmentation, more positive target entities are introduced after fetching the additional knowledge. However, for some negative target entities that fulfill positive attributes, the retrieved knowledge may contain the introduction about its positive attributes, conversely misleading the model to overlook its negative attributes. For example, for the negative target entity “Xinxiang” in the second column, it also belongs to Henan province, i.e., it satisfies the requirement of positive attributes. Therefore, after concatenating the retrieved entity introduction “Xinxiang is a prefecture-level city in northern Henan province, China.”, it is incorrectly introduced instead.
@@
Table 13. The prompt used to select the entities that are consistent with the attributes of $S^{pos}$ ($S^{neg}$) in the top-$T$ entities of $L_{0}$ to construct $L_{pos}$ ($L_{neg}$).
I have a task that involves classifying candidate entities based on their alignment with a seed entity set. The seed entities are grouped together because they share certain attributes, referred to as seed attributes. I will provide a list of seed entities along with their seed attributes. Additionally, I have a list of candidate entities that are similar to the seed entities but may not necessarily share the same seed attributes. I need you to identify the seed attributes and use them to classify each candidate entity into one of two categories: 1) consistent with the seed entity set in terms of seed attributes, or 0) inconsistent with the seed entity set in terms of seed attributes. For the given N candidate entities, please output N values, each being 1 or 0, indicating whether each candidate is consistent (1) or inconsistent (0) with the seed entity set based on the seed attributes.
Input:
Seed entities: [Mark Twain, Ernest Hemingway, F. Scott Fitzgerald]
Candidate entities: [J.K. Rowling, Stephen King, Agatha Christie, John Steinbeck, Harper Lee, Charles Dickens, Virginia Woolf], total 7 entities
Output:
”result”: [0,1,0,1,1,0,0]
Input:
Seed entities: [Golden Retriever, German Shepherd, Labrador Retriever]
Candidate entities: [Bengal Tiger, Beagle, Siberian Husky, African Elephant, Pug], total 5 entities
Output:
”result”: [0,1,1,0,1]
Input:
Seed entities: [{Entity1}, {Entity2}, {Entity3}]
Candidate entities: [{Entity1’}, {Entity2’}, {Entity3’}, …], total {} entities
Output:
Table 14. The prompt used to generate entities that are semantically similar to 3 given entities.
iron, copper, aluminum and zinc.
math, physics, chemistry and biology.
{Entity1}, {Entity2}, {Entity3} and ______
Table 15. The prompt used to generate a class name that covers the given entities.
Generate a class name that accurately represents the following entities. This class name should encompass all the given entities and reflect their shared characteristics.
Examples:
[[[[Tiger, Lion, Cheetah]]]] $\rightarrow$ Big Cats
[[[[Shakespeare, Tolstoy, Hemingway]]]] $\rightarrow$ Famous Authors
[[[[Mercury, Venus, Mars]]]] $\rightarrow$ Planets in the Solar System
[…]

### v3

@@ VI-D1 Analysis of Contrastive Learning
VI-D2 Analysis of Chain-of-thought Reasoning
TABLE VIII: Comparison experiments on chain-of-thought reasoning. The impact of reasoning depth and reasoning precision on performance is investigated. GT: Ground-truth. Gen: Generated. CN: Class name. Pos/Neg: Positive/Negative attribute information.
@@ VI-D2 Analysis of Chain-of-thought Reasoning
Experiments in Table VIII examine the chain-of-thought with varying reasoning depths and assess the impact of reasoning precision on model performance. From the results, we observe that: (1) Deeper reasoning proves to be more advantageous for Ultra-ESE, as it facilitates a more explicit understanding of the implicit ultra-fine-grained semantics of seed entities.
(2) Surprisingly, the use of manually labeled ground-truth class names is not superior to those derived through reasoning with the LLaMA-7B. The case study of the generated class names reveals that LLaMA-7B can produce class names encapsulating positive attribute information based on the given positive seed entities. For instance, when given entities about U.S. airports with the “location” attribute, GenExpan accurately infers more specific semantic classes that reflect positive attribute information, e.g., “Airports in Michigan”. This suggests that reasoning results are not necessarily inferior to manually annotated “ground-truth” labels, as chain-of-thought reasoning induces LLMs to attend to details that human ignores.
@@
Figure 7: Case studies of GenExpan and the chain-of-thought reasoning technique. $+$ and - - - stand for positive and negative target entities, respectively. !!! represents irrelevant entities belonging to the same fine-grained semantic class as the seed entities.
VI-E Case Study
@@
Figure 7 presents some case studies of GenExpan and the chain-of-thought reasoning technique. Overall, it is evident that GenExpan predominantly avoids expanding entities that do not belong to the same semantic class as the positive and negative seed entities. This emphasizes the relative tractability of traditional ESE tasks compared to the significant challenges posed by Ultra-ESE for existing LLMs, where the model cannot exclude the expansion of negative target entities and irrelevant entities (belonging to the same fine-grained semantic class).
@@
VII Conclusion
In conclusion, our study addresses the challenge of ultra-fine-grained Entity Set Expansion (Ultra-ESE) by introducing negative seed entities alongside positive ones, mitigating ambiguity and facilitating the expression of “unwanted” attributes. The creation of UltraWiki, a tailored dataset for Ultra-ESE, enables rigorous evaluation and future research. Through RetExpan and GenExpan frameworks, we demonstrate the effectiveness of large language models in Ultra-ESE from retrieval-based and generation-based perspectives. Our proposed strategies enhance models’ comprehension of ultra-fine-grained entity semantics. While our findings are promising, they also highlight the substantial scope for improvement in Ultra-ESE, advancing the field and paving the way for future exploration.
Acknowledgement
This research project is supported by National Natural Science Foundation of China (Grant No.62276154), Research Center for ComputerNetwork (Shenzhen) Ministry of Education, the Natural Science Foundation of Guangdong Province (Grant No.2023A1515012914 and 440300241033100801770), Basic Research Fund of Shenzhen City (Grant No.JCYJ20210324120012033, JCYJ20240813112009013 and GJHZ20240218113603006), the Major Key Projectof PCL for Experiments and Applications (PCL2022A05 and PCL2023A09), NSF under Grants III-2106758, and POSE-2346158.
Xuming Hu was supported by the Guangdong Provincial Department of Education Project (Grant No.2024KQNCX028); CAAI-Ant Group Research Fund; Scientific Research Projects for the Higher-educational Institutions (Grant No.2024312096), Education Bureau of Guangzhou Municipality; Guangzhou-HKUST(GZ) Joint Funding Program (Grant No.2025A03J3957), Education Bureau of Guangzhou Municipality.